### v2

@@ 3.3 JESTR vs explicit-construction models
JESTR is compared against ESP and MIST (Table 2A). For the NPLIB1 dataset, JESTR outperforms ESP and MIST on all reported ranks. For rank@1, JESTR outperforms ESP by 93.2%, and MIST by 64.1%. Further, JESTR achieves 95.8% rank@20, while the maximum performance of ESP and MIST is at 60.2% and 82.1%, respectively. We plot the detailed ranks for the NPLIB1 dataset (Figure 4A). At all ranks, JESTR provides superior performance to both ESP and MIST. For the NIST2020 dataset, JESTR outperforms all other models. Specifically, JESTR outperforms MIST at rank@1 by 83.8%. Detailed ranks for NIST2020 is provided in Figure S2A in the Supplementary Information. For the MoNA dataset, JESTR consistently outperforms ESP. MIST only outperforms JESTR at rank@1 by 17.6%, but not on rank@5, rank@20, or any other rank, as provided in Figure S3A in the Supplementary Information. Across all datasets, on avarage for ranks@[1-5], JESTR outperforms ESP by 71.6% and MIST by 23.6%.
@@
Examining the overall performance of JESTR over the three datasets, we note that JESTR’s performance was worse on the MoNA dataset when compared to NPLIB1 and NIST2020. We suspect that JESTR’s performance on MoNA was low for two reasons. First, MoNA has the fewest number of molecules, thus providing the lowest molecular diversity among the three datasets. Second, MoNA data is uploaded by users and may not have undergone consistent curation efforts. The NIST2020 dataset is well curated; however, it has the highest ratio of spectra per molecule, an average of 13.25 spectra per molecule, versus 1.13 and 5.28 spectra per molecule for NPLIB1 and MoNA, respectively. As such, we suspect that JESTR finds it hard to place all the spectra embeddings closer to the molecule in the joint space for NIST2020 and MoNA. A combination of high molecule to spectra ratio, lower diversity of molecules and inconsistent spectra curation makes JESTR perform the lowest on MoNA. MIST, with additional information in the form of subformulae annotation on peaks, does a better job at distinguishing among various spectra of the same molecule for MoNA, thereby attaining a better rank@1 score on this dataset. However, MIST looses its advantage over JESTR starting with rank@2.
@@ 3.5 Ablation study - removing regularization
To assess the value of regularization, the model was retrained without . Regularizing the training loss with molecular candidates improves rank@1 by 11.4%, 6.0% and 37.1% on the NPLIB1, NIST2020 and MoNA datasets, respectively. Improvements using regularization is evident at almost all ranks and all datasets ((Table 2B). For rank@20 on NPLIB1, the results drop by 0.3% since the rank@20 result for NPLIB1 is high even without regularization, where even a small change in a few targets changes the result slightly.
@@ 4 Conclusion
JESTR offers a novel implicit annotation paradigm that avoids the explicit generation of spectra, fingerprints, or molecular structures. As molecules and spectra are views of the same object, embedding these views in a joint embedding space using contrastive learning provides a performance advantage. On NPLIB1, JESTR outperforms ESP by 88.9% and MIST by 41.1% on ranks@[1-5]. On NIST2020, JESTR outperforms ESP by 68.5% and MIST by 25.8% on average on rank@[1-5]. On MoNA, JESTR outperforms ESP by 57.5% and MIST by 4.0% on average on rank@[1-5]. Analysis of JESTR’s performance on the three datasets reveals that dataset diversity, quality, and spectra-to-molecule ratios impact performance. Further, our results showed enhance value in utilizing candidate molecules during training for regularization, improving performance an average of 11.4% for all three datasets. The overall JESTR results are promising and vouch for the potential of implicit annotation approaches. We expect to attain further improvements by utilizing additional knowledge in the form of subformulae annotation on spectral peaks and by utilizing enhanced molecular and spectral encoders.
@@
Research reported in this publication was supported by the National Institute of General Medical Sciences of the National Institutes of Health under award number R35GM148219.
Supplementary Information
S1 Batching strategy for contrastive learning and regularization
During contrastive learning, each batch of size $k$ contains $k$ target molecules. In each training epoch, the spectrum corresponding to any given target molecule is chosen sequentially from the list of paired spectra. The batch size $k$ was set to 32. The value of the temperature hyper-parameter $\tau$ was set to 0.05. For all experiments, contrastive learning was stopped at 1000 epochs, where it was observed that the contrastive loss stopped reducing significantly. The batch size for the candidates, $k_{aug}$, was set to 32.
During regularization, for each training molecule in the batch, a set of most similar candidates for that molecule are used to calculate the regularization loss. This is done sequentially for every mini-batch in the epoch. A diagram outlining our batching strategy is shown in Figure S1.
Figure S1: The contrastive loss is calculated over a batch of molecules at a time. For each molecule in the batch, candidates are selected from the candidate list sorted by similarity. If a particular molecule has a lesser number of candidates than the batch size, the candidates for that molecule will be repeated sequentially. The candidates thus selected are used to calculate the regularization loss. The curly braces in the vertical direction show the contrasting batching for two batches, while the curly braces in the horizontal direction show regularization batches for two batches
S2 Implementation details for JESTR
JESTR was trained on NVIDIA A100 GPU with 40GB of graphics RAM and 256GB of CPU RAM. Adam (Kingma and Ba, 2014) was used as the optimizer. We used grid search over the ranges of the tuned parameters. The values of the parameters that achieved the best performance were selected and used to train and test the model for all datasets (Table S1).
Table S1: Tuning of various hyperparameters for JESTR model using grid search.
Hyperparameter Values searched Final value
Contrastive Learning Rate 5e-6, 5e-5, 5e-4 5e-4
Contrastive Batch Size 32, 64, 128 32
Early Stopping Epochs 20, 40, 80 80
Regularization loss weight 0.1, 0.3, 0.5 0.1
% epochs used for regularization 3, 10, 20 3
We also looked at the training and inference times of all the models (Table S2). The runtimes were measured on a Linux machine with 48 CPU cores with 196GB RAM and 6 nVidia A5000 GPUs with 24GB RAM. JESTR uses a GNN encoder for the molecules and an MLP encoder for the spectra. MIST uses MLP and transformers as its encoders for spectra, while ESP uses MLP and GNN as encoders for spectra and molecules, respectively. From a computational complexity of the model architecture, the three models are similar. The training time depends largely on the amount of training each model undergoes.
Table S2: Training time and inference time for the 3 models on the NPLIB1 dataset. The runtimes were measured on the same machine. JESTR takes longer to train because the training is run for a larger number of epochs.
Model Training Time Epochs Inference Time
JESTR 6 hours 800 1.5 hours
MIST 1.2 hours 50 20 minutes
ESP 1.3 hours 100 1.2 hours
S3 JESTR ranks and target-candidate separation on NIST2020 and MoNA
We look at how JESTR separates target and candidate molecules for different datasets in Figures (S2 and S3)
Figure S2: Results. A. Rank@k results for JESTR on NIST2020, with and without regularization, MIST, and ESP MLP-PD. B. Distribution of cosine similarities of query spectra and target/candidate molecules in the NIST2020 test set with contrastive learning using JESTR.
Figure S3: Results. A. Rank@k results for JESTR on MoNA, with and without regularization, MIST, and ESP MLP-PD. B. Distribution of cosine similarities of query spectra and target/candidate molecules in the MoNA test set with contrastive learning using JESTR.
S4 Regularization analysis for NIST2020 and MoNA
We study the impact of regularization on molecule ranking and on molecule embeddings for all datasets in Figures (S4 and S5).
Figure S4: Regularization analysis for JESTR for NIST2020. A. Regularization improves rank@k by significantly placing more targets at rank 1. B. Distribution on Tanimoto similarities on the ECFP fingerprints between target and candidates in the training set. C. Distribution on cosine similarities, with and without regularization, of the target and candidates within the test set.
Figure S5: Regularization analysis for JESTR for MoNA. A. Regularization improves rank@k by significantly placing more targets at rank 1. B. Distribution on Tanimoto similarities on the ECFP fingerprints between target and candidates in the training set. C. Distribution on cosine similarities, with and without regularization, of the target and candidates within the test set.

### v3

@@ 3.3 JESTR vs explicit-construction models
JESTR is compared against ESP and MIST (Table 2). For the NPLIB1 dataset, JESTR outperforms ESP and MIST on all reported ranks. For rank@1, JESTR outperforms ESP by 93.2%, and MIST by 64.1%. Further, JESTR achieves 95.8% rank@20, while the maximum performance of ESP and MIST is at 60.2% and 82.1%, respectively. We plot the detailed ranks for the NPLIB1 dataset (Figure 4A). At all ranks, JESTR provides superior performance to both ESP and MIST. For the NIST2020 dataset, JESTR outperforms all other models. Specifically, JESTR outperforms MIST at rank@1 by 83.8%. Detailed ranks for NIST2020 is provided in Figure S2A in the Supplementary Information. For the MoNA dataset, JESTR consistently outperforms ESP. MIST only outperforms JESTR at rank@1 by 17.6%, but not on rank@5, rank@20, or any other rank, as provided in Figure S3A in the Supplementary Information. On the MassSpecGym dataset, JESTR achieves 15.62% rank@1 using candidates by mass, a 45.9% improvement over ESP and 6.7% improvement over MIST. Using candidates by formula, JESTR achieves 11.85% rank@1, surpassing ESP by 7.2% and MIST by 23.8%.
@@
Examining the overall performance of JESTR against ESP and MIST, over the four datasets, on average for rank@[1-20], JESTR outperforms ESP by 55.5% and MIST by 56.6%. JESTR’s performance was worse on the MoNA dataset when compared to NPLIB1 and NIST2020. We suspect that JESTR’s performance on MoNA was low for two reasons. First, MoNA has the fewest number of molecules, thus providing the lowest molecular diversity among the three datasets. Second, MoNA data is uploaded by users and may not have undergone consistent curation efforts. The NIST2020 dataset is well curated; however, it has the highest ratio of spectra per molecule, an average of 13.25 spectra per molecule, versus 1.13 and 5.28 spectra per molecule for NPLIB1 and MoNA, respectively. As such, we suspect that JESTR finds it hard to place all the spectra embeddings closer to the molecule in the joint space for NIST2020 and MoNA. A combination of high molecule to spectra ratio, lower diversity of molecules and inconsistent spectra curation makes JESTR perform the lowest on MoNA. MIST, with additional information in the form of subformulae annotation on peaks, does a better job at distinguishing among various spectra of the same molecule for MoNA, thereby attaining a better rank@1 score on this dataset. However, MIST loses its advantage over JESTR starting with rank@2.
As SIRIUS and CFM-ID are widely used within the community, we benchmark JESTR’s performance against their publicly available pretrained models. To ensure a fair evaluation, we focus our evaluation on the MassSpecGym benchmark dataset using candidates by formula. Among the datasets, MassSpecGym offers the most equitable basis for comparison, as the distribution of MCES distances between the pretrained molecules in SIRIUS and CFM-ID and the test molecules in MassSpecGym closely matches that of the dataset’s train/test split (Figures S6-S8). However, because some test molecules were included in the training of SIRIUS and CFM-ID, these models may benefit from an advantage not available to JESTR. To address this, we report results on both the full test set and a filtered subset excluding any molecules seen during baseline model pretraining.
For SIRIUS, we create a custom database containing all candidates in the test dataset. We use the GUI version 6.1.0 and ensure the correct adduct and instrument setting for each spectrum is used. Out of 17,556 test spectra, SIRIUS assigned at least one structure to 7,766 spectra, with 3,342 of those with the target molecule ranked in position 1, that is a 19.04% rank@1 performance compared to JESTR at 11.85%. Excluding test spectra whose molecules were in SIRIUS’s training set yields 1,281 spectra correctly annotated out of 12,314 spectra, a 10.41% rank@1 vs 13.65% for JESTR. This result translates to a 31.12% improvement over SIRIUS, indicating that JESTR generalizes better than SIRIUS on unseen molecules.
For CFM-ID, per the authors’ recommendation, we combine 3 spectra of low (<20eV), medium ($\geq$ 20eV and <40eV), and high ($\geq$ 40eV) collision energies per test molecule. Spectra are chosen randomly within the threshold, and we skip a spectrum of a specific collision energy threshold if there is not a spectrum whose collision energy falls within the threshold. Since JESTR is not designed to rank based on a set of spectra, we report the averages for best (lowest rank) and median performance across the chosen spectra. Out of 3,170 merged spectra, CFM-ID correctly annotates 62 spectra, achieving 1.96% rank@1 vs 14.10% rank@1 based on the best performance and 6.56% rank@1 based on the median rank for JESTR. Removing test spectra whose molecules were in the training set yields 60 spectra correctly annotated out of 3,004 spectra, a 2.00% rank@1 vs 14.61% rank@1 based on the best rank and 6.76% rank@1 based on the median rank for JESTR. Based on the median rank, JESTR provides a 238.00% improvement over CMF-ID.
The MassSpecGym benchmark dataset is notably more difficult than other datasets for all models, as molecules in the test set drastically differ from those in the training set. Nevertheless, JESTR’s superior performance against explicit-construction models highlights the effectiveness of an implicit model using the joint embedding space for ranking candidate molecules.
3.4 JESTR vs implicit models
Similar to JESTR, CMSSP learns a joint embedding space for spectra annotation using contrastive learning. We report CMSSP’s performance when training on individual datasets and using the released model weights, as training CMSSP on individual datasets shows low performance. JESTR significantly outperforms CMSSP across all datasets except for rank@1 on the NPLIB1 dataset against the pretrained CMSSP model (Table 2). The strong performance of CMSSP likely results from similarities, or potentially data leakage, between our NPLIB1 test set and CMSSP training data.
We suspect CMSSP performed poorly when it was retrained because of its large model architecture, requiring the training of a large number of parameters. Specifically, CMSSP uses a GNN-based encoder for the molecule together with an MLP for the molecular fingerprint while using a transformer-based encoder with 8 attention heads for the spectra. Two other differences between CMSSP and JESTR is the loss function and the use of the dot product vs cosine similarity to evaluate embedding similarities. We evaluate the impact of these differences on the NPLIB1 dataset (Table 3). We first train the JESTR architecture with CMSSP’s cross entropy (CE) loss and the dot product. Performance drops substantially when swapping the CMSSP architecture for JESTR’s architecture. Based on performance, the CMSSP architecture is more suited for using CE loss and the dot-product when compared to the JESTR architecture. Maintaining the CE loss, we then evaluate the effect of swapping the cosine ranking to replace the dot-product ranking. Using cosine similarity for ranking surpasses using the dot product and improves rank@5 and rank@20 performance over the CMSSP model by 22.5% and 36.8%, respectively, even though it was trained using the dot product. Cosine similarity explicitly normalizes the embedding magnitudes, yielding a competitive edge: similarity comparisons rely purely on embedding direction rather than magnitude. Embeddings with artificially large magnitudes are prevented from dominating similarity scores. InfoNCE with temperature-scaled cosine similarity is more suitable for a contrastive learning framework, as the normalization from cosine similarity reduces the influence of magnitude differences and the temperature scaling allows for fine-grained control over the distribution of similarities. The combination of a simple architecture and InfoNCE with temperature-scaled cosine similarity provides JESTR a superior performance compared to CMSSP.
Table 3: Ranking result on the NPLIB1 dataset, contrasting the CMSSP architecture against the JESTR architecture with varying losses: cross entropy (CE) and InfoNCE losses, and ranking with dot product (DP) and cosine similarity.
CMSSP Architecture JESTR Architecture
CE loss CE loss CE loss InfoNCE loss
Rank (%) DP ranking DP ranking cosine ranking cosine ranking
rank@1 13.92 3.20 10.84 45.76
rank@5 32.97 16.38 40.39 81.53
rank@20 49.08 37.68 67.12 95.77
@@ 3.6 Ablation study - removing regularization
To assess the value of regularization, the model was retrained without. Regularizing the training loss with molecular candidates improves rank@1 by 11.4%, 6.0% and 37.1% on the NPLIB1, NIST2020 and MoNA datasets, respectively. Improvements using regularization are evident at almost all ranks and all datasets (Table 2B). For rank@20 on NPLIB1, the results drop by 0.3% since the rank@20 result for NPLIB1 is high even without regularization, where even a small change in a few targets changes the result slightly. On the MassSpecGym dataset, regularization did not consistently boost model performance. Specifically, we observe a slight decrease in performance across all rank@k’s when evaluating candidates by mass, and similarly, a modest reduction in rank@5 performance when using candidates by formula. We suspect that the challenging split within this dataset limits the generalizability benefits provided by regularization.
@@ 4 Conclusion
JESTR offers a novel implicit annotation paradigm that avoids the explicit generation of spectra, fingerprints, or molecular structures. As molecules and spectra are views of the same object, embedding these views in a joint embedding space using contrastive learning provides a performance advantage. Evaluation across diverse datasets, such as NPLIB1, NIST2020, MoNA, and MassSpecGym, demonstrates clear performance advantages over current explicit annotation techniques, highlighting the robustness of implicit models using the joint embedding approach. Moreover, investigation on the contrastive loss used by JESTR reveals that InfoNCE with temperature-scaled cosine similarity is advantageous for contrasting views and learning to distinguish matching pairs from non-matching pairs. Analysis of JESTR’s performance on these datasets reveals that dataset diversity, quality, and spectra-to-molecule ratios impact performance. Further, our results showed enhanced value in utilizing candidate molecules during training for regularization, improving performance by an average of 5.72% for all datasets. Nonetheless, the minimal improvement in performance with regularization on the MassSpecGym benchmark suggests that the effectiveness of regularization may vary depending on data splits and generalization difficulty between the train and test set. The overall JESTR results are promising and vouch for the potential of implicit annotation approaches. We expect to attain further improvements by utilizing additional knowledge in the form of subformulae annotation on spectral peaks and by utilizing enhanced molecular and spectral encoders.
@@
Research reported in this publication was supported by the National Intitute of General Medical Sciences of the National Institutes of Health under award number R35GM148219. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.